OpenStudy (anonymous):

Find the zeros of the polynomial function, and state the multiplicity of each.

OpenStudy (anonymous):

y = 4(x + 7)^2(x - 7)^3

OpenStudy (anonymous):

These are the options: A. 4, multiplicity 1; -7, multiplicity 3; 7, multiplicity 3 B. -7, multiplicity 3; 7, multiplicity 2 C. 4, multiplicity 1; 7, multiplicity 1; -7, multiplicity 1 D. -7, multiplicity 2; 7, multiplicity 3

hartnn (hartnn):

to find zeros, u put y=0 so u have 4(x + 7)^2(x - 7)^3 =0 that means (x + 7)^2=0 OR (x - 7)^3=0

OpenStudy (anonymous):

So how do I find the zeros from those equations?

hartnn (hartnn):

when (x-a)=0, a is a zero so when (x+7)=0 what will be the zero ?

OpenStudy (anonymous):

not that easy yes to find zeros you have to put y=0 and then use wavy curve method to solve the zero

OpenStudy (anonymous):

I am so confused.

OpenStudy (anonymous):

hartnn, is it -7?

hartnn (hartnn):

yes! and since it has power = 2 its multiplicity = 2

hartnn (hartnn):

now u also have (x - 7)^3 =0 so what is a zro here ?

OpenStudy (anonymous):

The zero is 7, and the multiplicity is 3. Is that right?

hartnn (hartnn):

absolutely correct!

OpenStudy (anonymous):

So the answer is B?

hartnn (hartnn):

7, and the multiplicity is 3 so its D.

OpenStudy (anonymous):

Oh, I see, whoops. Thanks so much! :)

hartnn (hartnn):

welcome :)